3.6gm of an ideal gas  was injected into a bulb of internal volume of 8.221L at pressure Patm And temp Tk . The bulb was then placed in a themostat maintained at (T+15)k. .6 gm of the gas was let off keep the original pressure . find Pand T if mol weight of gas is 36

Grade:12th pass

1 Answers

Prakhar singh
2 years ago
T =75 k
P=0.075
Use formula n1t1=n2t2 
And get the value of t and then put it into 
PV=nRT 
N1=1/10 moles
N2 = 1/12moles 
T1= T k
T2= (T+15)k 
 
n1t1=n2t2
T=75k 
PV=nRT
P= 0.0821*75/82.1= 0.075 atm